OpenStudy (anonymous):

A system of equations is shown below. 8x + 5y = 9 3x + 2y = 4 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

Can someone please help me? Im on a tight schedule and I need the answers

OpenStudy (anonymous):

Please help???!!!

OpenStudy (anonymous):

8x+5y=9 3x+2y=4 8x+5y=9(x2) 3x+2y=4(-5) 16x+10y=18 -15x-10y=-20 x=-2

OpenStudy (anonymous):

Substitution (8x-2)+5y=9 -16+2y=4 2y=4+16 2y=20 y=10